OpenStudy (anonymous):

Choose the correct simplification of the expression (2x7y)2(y5)3. PLEASE HELP!!!

OpenStudy (anonymous):

Do you need an explanation or just the answer?

OpenStudy (anonymous):

4x7y10 4x9y10 4x14y16 4x14y17

OpenStudy (anonymous):

both please!!

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

[(2x^7y)^2] (y^5)^3 (4x^14y^2)(y^15) 4x^14y^17

OpenStudy (anonymous):

can you help with one more @choopers ??

OpenStudy (anonymous):

Choose the correct simplification of the expression a^7b^8/a^4b^4

OpenStudy (anonymous):

just subtract the exponent. *a - a = 7 - 4 = 3 so, a^3 *b - b = 8 - 4 = 4 so, b^4 the answer is: a^3 b^4

OpenStudy (anonymous):

Good job choopers i would have helped but I was helping someone else.

OpenStudy (anonymous):

thank you so so much i really appreciate it im starting to understand! @choopers

OpenStudy (anonymous):

@Puahina thanks. :) @mathhelpplease11 no problem. :)